# AdaSCALE: Adaptive Scaling for OOD Detection

### v1

@@ Abstract
The ability of the deep learning model to recognize when a sample falls outside its learned distribution is critical for safe and reliable deployment. Recent state-of-the-art out-of-distribution (OOD) detection methods leverage activation shaping to improve the separation between in-distribution (ID) and OOD inputs. These approaches resort to sample-specific scaling but apply a static percentile threshold across all samples regardless of their nature, resulting in suboptimal ID-OOD separability. In this work, we propose AdaSCALE, an adaptive scaling procedure that dynamically adjusts the percentile threshold based on a sample’s estimated OOD likelihood. This estimation leverages our key observation: OOD samples exhibit significantly more pronounced activation shifts at high-magnitude activations under minor perturbation compared to ID samples. AdaSCALE enables stronger scaling for likely ID samples and weaker scaling for likely OOD samples, yielding highly separable energy scores. Our approach achieves state-of-the-art OOD detection performance, outperforming the latest rival OptFS by 14.94% in near-OOD and 21.67% in far-OOD datasets in average FPR@95 metric on the ImageNet-1k benchmark across eight diverse architectures. The code is available at: https://github.com/sudarshanregmi/AdaSCALE/
@@ 1 Introduction
The reliable deployment of deep learning models hinges on their ability to handle previously unseen inputs, a task commonly known as OOD detection. One critical application is in medical diagnosis, where a model trained on common diseases should be able to flag inputs representing unknown conditions as potential outliers, requiring further review by clinicians. OOD detection primarily involves identifying semantic shifts, with robustness to covariate shifts being a highly desirable characteristic [58, 2]. As modern deep learning models scale in both data and parameter counts, effective OOD detection within large-scale settings is critical. Given the difficulties of iterating on large models, post-hoc approaches that preserve ID accuracy are generally preferred.
@@
A variety of post-hoc approaches have emerged, broadly categorized by where they operate. One class of methods focuses on computing OOD scores directly in the output space [16, 30, 33, 20, 7, 64], while another operates in the activation space [27, 48, 44, 40]. Finally, a more recent line of research also explores a hybrid approach [50, 24], combining information from both spaces. The efficacy of many high-performing methods relies on either accurate computation of ID statistics [47, 25, 55, 46, 26] or retention of training data statistics [48, 40]. However, as retaining full access to training data becomes increasingly impractical in large-scale settings, methods that operate effectively with minimal ID samples without performance degradation are particularly valuable for practical applications.
@@
Alleviating the dependence on ID training data/statistics, recent state-of-the-art post-hoc approaches center around the concept of “fixed scaling.” ASH [7] prunes and scales activations on a per-sample basis. SCALE [54], the direct successor of ASH, critiques pruning and focuses purely on scaling, which improves OOD detection without accuracy degradation. LTS [8] extends this concept by directly scaling logits instead, using post-ReLU activations. These methods leverage a key insight: scaling based on the relative strength of a sample’s top-k activations (with respect to the entire activations) produces highly separable ID-OOD energy scores. However, although such approaches provide sample-specific scaling factors, the scaling mechanism remains uniform across all samples as the percentile threshold and thereby is fixed, as shown in Figure 1. This static approach is inherently limiting for optimal ID-OOD separation while also failing to leverage even minimal ID data, which could be reasonably practical in most deployment scenarios.
@@
We hypothesize that designing an adaptive scaling procedure based on each sample’s predetermined OOD likelihood offers greater control for enhancing ID-OOD separability. Specifically, this mechanism should assign smaller scaling factors for samples with high OOD likelihood to yield lower energy scores and larger scaling factors for probable ID samples to yield higher energy scores. To achieve this, we propose a heuristic for predetermining OOD likelihood based on a key observation in activation space: minor perturbations applied to OOD samples induce significantly more pronounced shifts in their top-k activations compared to ID samples. Consequently, samples exhibiting substantial activation shifts are assigned lower scaling factors, while those with minimal shifts receive higher scaling factors. This adaptive scaling mechanism can be applied in either logit or activation space. Our method, AdaSCALE, achieves state-of-the-art performance, delivering significant improvements in OOD detection while requiring only minimal ID samples.
@@
We conduct an extensive evaluation across 8 architectures on ImageNet-1k and 2 architectures on CIFAR benchmarks, demonstrating the substantial effectiveness of AdaSCALE. For instance, AdaSCALE surpasses the average performance of the best-generalizing method, OptFS [64], by for near-OOD detection and for far-OOD detection in terms of FPR@95 / AUROC, on the ImageNet-1k benchmark across eight architectures. Furthermore, AdaSCALE outperforms the best-performing method, SCALE [54], when evaluated on the ResNet-50 architecture, achieving performance gains of for near-OOD and for far-OOD detection. Additionally, AdaSCALE consistently demonstrates superiority in full-spectrum OOD detection [58]. Our key contributions are summarized as follows:
@@
We reveal that OOD inputs exhibit significantly more pronounced shifts in top-k activations under minor perturbations compared to ID inputs.
Leveraging this observation, we propose a novel post-hoc OOD detection method utilizing adaptive scaling that attains state-of-the-art OOD detection.
@@
We demonstrate state-of-the-art generalization of AdaSCALE via extensive evaluations across many setups.
@@ 2 Related Works
Early research on OOD detection primarily focused on designing scoring functions based on logit information [16, 30, 33, 20, 34]. While these methods leveraged logit-based scores, alternative approaches have explored gradient-based information, such as GradNorm [23], GradOrth [4], GAIA [5], and Greg-OOD [45]. Given the limited dimensionality of the logit space, which may not encapsulate sufficient information for OOD detection, subsequent studies have investigated activation-space-based methods. These approaches exploit the high-dimensional activations, leading to both parametric techniques such as MDS [27], MDS Ensemble [27], and RMDS [44], as well as non-parametric methods such as KNN-based OOD detection [48, 39]. Recent advancements have proposed hybrid methodologies that integrate parametric and non-parametric techniques to improve robustness. For instance, ComboOOD [40] combines these paradigms to enhance near-OOD detection performance. Similarly, VIM [50] employs a combination of logit-based and distance-based metrics. However, reliance of such approaches on ID statistics [46, 38, 59] can become a constraint, hindering scalability and practical deployment in real-world applications. To mitigate computational challenges for real-world deployment, recent methods, such as FDBD [32] and NCI [31], have focused on enhancing efficiency. Recent advances, such as WeiPer [14], explore class-direction perturbations, while NECO [1] examines connections to neural collapse phenomena.
Activation-shaping post-hoc methods.
@@
A seminal work in OOD detection, ReAct [47], identified abnormally high activation patterns in OOD samples and proposed clipping extreme activations. This approach has been further generalized by BFAct [25] and VRA [55], which extend activation clipping for enhanced effectiveness. Additionally, BATS [66] refines activation distributions by aligning them with their respective typical sets, while LAPS [15] enhances this strategy by incorporating channel-aware typical sets. Inspired by activation clipping, another line of research explores activation “scaling” as a means to improve OOD detection. ASH [7] introduces a method to compute a scaling factor as a function of the activation itself, pruning and rescaling activations to enhance the separation of energy scores between ID and OOD samples. However, this approach results in a slight degradation in ID classification accuracy. In response, SCALE [54] observes that pruning adversely affects performance and thus eliminates it, leading to improved OOD detection while preserving ID accuracy. SCALE currently represents the state-of-the-art method for ResNet-50-based OOD detection. Despite their efficacy, these activation-based methods exhibit limited generalization across diverse architectures. To address this issue, LTS [8] extends SCALE by computing scaling factors using post-ReLU activations and applying them directly to logits rather than activations. Our work builds on this line of work, introducing the adaptive scaling mechanism. ATS [26] argues that relying solely on final-layer activations may result in the loss of critical information beneficial for OOD detection and proposes to leverage intermediate-layer activations too. However, its efficacy is contingent upon the availability of a large number of training samples, whereas our approach attains state-of-the-art performance while utilizing a minimal number of ID samples. A newly proposed method OptFS [64] introduces a piecewise constant shaping function with the goal of generalization across diverse architectures in large-scale settings, while our work exhibits superior generalization extending to small-scale settings too.
Training methods.
@@
The training methods incorporate adjustments during training to enhance the ID-OOD differentiating characteristics. They either make architectural adjustments [6, 18, 22], apply enhanced data augmentations [53, 19, 21], or make simple training modifications [52, 42, 62]. More recent methods have adopted contrastive learning in the context of OOD detection [36, 43, 35, 67]. Moreover, some approaches also either utilize external real outliers [17, 60, 11, 65] or synthesize virtual outliers either in image space [10, 41, 51, 13, 63, 29, 3, 37] or in feature space [9, 49, 12, 28]. However, training strategies incur additional computational cost and potentially inferior performance to post-hoc methods in large-scale setups [57].
@@ 3 Preliminaries
Let denote the input space and denote the label space, where is the number of classes. We consider a multi-class classification setting where a classifier is trained on ID data drawn from an underlying joint distribution , where and . The ID training dataset is denoted as , where is the number of training samples and .
The classifier is composed of a feature extractor , and a classifier . The feature extractor maps an input to a feature vector , where and the classifier then maps this feature vector to a logit vector . We refer to individual dimensions of the feature vector as activations, denoted by for the -th dimension. The classifier is trained on to minimize the empirical risk:
where is a loss function, such as cross-entropy loss. During inference, the model may encounter data points drawn from a different distribution, denoted as , which is referred to as OOD data. The OOD detection problem aims to identify whether a given input is drawn from marginal distribution or from . Hence, the goal is to design a scoring function that assigns a scalar score to each input , reflecting its likelihood of being an OOD sample. A higher score typically indicates a higher probability of the input being OOD. A threshold is used to classify an input as either ID or OOD:
@@ 4 Method
In this section, we introduce AdaSCALE, a novel post-processing approach that dynamically adapts the scaling mechanism based on each sample’s estimated OOD likelihood. We first present our key empirical observations regarding activation behavior under minor perturbations, building upon insights from ReAct [47]. Next, we revisit and analyze the core principle underlying recent state-of-the-art approaches. Finally, we detail our proposed adaptive scaling mechanism that leverages these observations to achieve superior OOD detection performance.
@@ 4.1 Observations in Activation Space
A seminal work ReAct [47] demonstrated that OOD samples often induce abnormally high activations within neural networks. We extend this finding with an important observation: the positions of such high activations in OOD samples are relatively unstable under minor perturbations compared to ID samples. This instability provides a valuable signal for distinguishing OOD samples from ID samples. Below, we formalize this observation and our methodology.
@@ 4.1.1 Perturbation Mechanism
Let be an input image with input channels, height, and width. We denote channel value at position as . To identify channel values for perturbation, we employ pixel attribution that quantifies each input element’s influence on the model’s prediction. An attribution function, , assigns a score to each channel value, with lower absolute scores indicating less influence. We select of channel value indices with lowest absolute attribution scores, forming the set . We use a gradient-based attribution:
where is the predicted class index. To create a perturbed input, we select a subset containing of channel values to perturb. The perturbed image is obtained as:
where is perturbation magnitude.
@@ 4.1.2 Activation Shift as OOD Indicator
After obtaining the perturbed input , we compute its activation . We define the activation shift as the absolute element-wise difference between the original activation and the perturbed activation:
@@
Figure 3 illustrates the key insight of our approach: activation shift at extreme (high-magnitude) activations is consistently more pronounced in OOD samples compared to ID samples. This behavior can be understood intuitively: ID samples activate network features in a stable, predictable manner reflecting learned patterns, while OOD samples trigger less stable, more arbitrary high activations that shift significantly under perturbation. Based on this observation, we propose using activation shift at the top- highest activations as a metric to estimate OOD likelihood of a sample:
where returns the indices of the highest values in .
@@
As evidenced by ratio shown in Figure 4, the statistic generally assigns higher values to diverse OOD samples than ID samples. However, the high variance of metric (Figure 3) suggests the possibility of overoptimistic estimations. To address this issue, we introduce a correction term that exhibits an opposing behavior: it tends to be higher for ID samples than for OOD samples. Figure 5 shows that the perturbed activations of ID samples tend to be higher than those of OOD samples, especially in high-activation regions. We leverage this complementary signal by defining:
where is a hyperparameter denoting the number of considered activations. We refine our OOD quantification by combining both metrics, weighted by a hyperparameter :
If be the set of values on ID validation samples, we could transform any into a normalized probability scale by constructing empirical cumulative distribution function (eCDF) derived from . The eCDF, denoted as , can be defined as:
@@
where is the indicator function. A higher value of indicates a higher likelihood of the sample being OOD. Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table 7).
4.2 Revisiting Static Scaling Mechanisms
Recent state-of-the-art post-processing methods for OOD detection, including SCALE [54], ASH [7], and LTS [8], operate by scaling either the activations or logits. The scaling factor is computed as:
where denotes the percentile of all elements in activation . While this approach yields sample-specific scaling factors, it imposes a critical constraint: the percentile threshold is static and identical across all test samples, regardless of the nature of samples. We argue that this static nature limits the effectiveness of the scaling procedure and prevents optimal ID-OOD separability.

### v2

@@ Abstract
The ability of the deep learning model to recognize when a sample falls outside its learned distribution is critical for safe and reliable deployment. Recent state-of-the-art out-of-distribution (OOD) detection methods leverage activation shaping to improve the separation between in-distribution (ID) and OOD inputs. These approaches resort to sample-specific scaling but apply a static percentile threshold across all samples regardless of their nature, resulting in suboptimal ID-OOD separability. In this work, we propose AdaSCALE, an adaptive scaling procedure that dynamically adjusts the percentile threshold based on a sample’s estimated OODness. This estimation leverages our key observation: OOD samples exhibit significantly more pronounced activation shifts at high-magnitude activations under minor perturbation compared to ID samples. AdaSCALE enables stronger scaling for likely ID samples and weaker scaling for likely OOD samples, yielding highly separable energy scores. Our approach achieves state-of-the-art OOD detection performance, outperforming the latest rival OptFS by 14.94% in near-OOD and 21.67% in far-OOD datasets in average FPR@95 metric on the ImageNet-1k benchmark across eight diverse architectures. The code is available at: https://github.com/sudarshanregmi/AdaSCALE/
@@ 1 Introduction
The reliable deployment of deep learning models hinges on their ability to handle previously unseen inputs, a task commonly known as OOD detection. A critical application is in medical diagnosis, where a model trained on common diseases should be able to flag inputs representing unknown conditions as potential outliers, requiring further review by clinicians. OOD detection primarily involves identifying semantic shifts, with robustness to covariate shifts being a highly desirable characteristic (Yang et al., 2023; Baek et al., 2024). As deep learning models scale in both data and parameter counts, effective OOD detection within large-scale settings is critical. Given the difficulties of iterating on large models, post-hoc approaches preserving ID accuracy are generally preferred.
@@
A variety of post-hoc approaches have emerged, broadly categorized by where they operate. One class of methods focuses on computing OOD scores directly in the output space (Hendrycks & Gimpel, 2017; Liang et al., 2018; Liu et al., 2020; Hendrycks et al., 2022a; Djurisic et al., 2023; Zhao et al., 2024), while another operates in the activation space (Lee et al., 2018; Sun et al., 2022; Ren et al., 2021; Rajasekaran et al., 2024). Finally, a more recent line of research also explores a hybrid approach (Wang et al., 2022; Kim et al., 2024), combining information from both spaces. The efficacy of many high-performing methods relies on either accurate computation of ID statistics (Sun et al., 2021; Kong & Li, 2022; Xu et al., 2023; Sun & Li, 2022; Krumpl et al., 2024) or retention of training data statistics (Sun et al., 2022; Rajasekaran et al., 2024). However, as retaining full access to training data becomes increasingly impractical in large-scale settings, methods that operate effectively with minimal ID samples are particularly valuable for practical applications.
@@
Alleviating the dependence on ID training data/statistics, recent state-of-the-art post-hoc approaches center around the concept of “fixed scaling." ASH (Djurisic et al., 2023) prunes and scales activations on a per-sample basis. SCALE (Xu et al., 2024), the direct successor of ASH, critiques pruning and focuses purely on scaling, which improves OOD detection performance without accuracy degradation. LTS (Djurisic et al., 2024) extends this concept by directly scaling logits instead, using post-ReLU activations. These methods leverage a key insight: scaling based on relative strength of a sample’s top-k activations (with respect to entire activations) yields highly separable ID-OOD energy scores. However, although such approaches provide sample-specific scaling factors, the scaling mechanism remains uniform across all samples as the percentile threshold and thereby is fixed, as shown in Figure˜1. This static approach is inherently limiting for optimal ID-OOD separation while also failing to leverage even minimal ID data, which could be reasonably practical in most deployment scenarios.
@@
We hypothesize that designing an adaptive scaling procedure based on each sample’s estimated OODness offers greater control for enhancing ID-OOD separability. Specifically, this mechanism should assign smaller scaling factors for samples with high OODness to yield lower energy scores and larger scaling factors for probable ID samples to yield higher energy scores. To achieve this, we propose a heuristic for estimating OODness based on a key observation in activation space: minor perturbations applied to OOD samples induce significantly more pronounced shifts in their top-k activations compared to ID samples. Consequently, samples exhibiting substantial activation shifts are assigned lower scaling factors, while those with minimal shifts receive higher scaling factors. This adaptive scaling mechanism can be applied in either logit or activation space. Our method, AdaSCALE, achieves state-of-the-art performance, delivering significant improvements in OOD detection while requiring only minimal ID samples.
@@
We conduct an extensive evaluation across 8 architectures on ImageNet-1k and 2 architectures on CIFAR benchmarks, demonstrating the substantial effectiveness of AdaSCALE. For instance, AdaSCALE surpasses the average performance of the best-generalizing method, OptFS (Zhao et al., 2024), by for near-OOD detection and for far-OOD detection in terms of FPR@95 / AUROC, on the ImageNet-1k benchmark across eight architectures. Furthermore, AdaSCALE outperforms the best-performing method, SCALE (Xu et al., 2024), when evaluated on the ResNet-50 architecture, achieving performance gains of for near-OOD and for far-OOD detection. Additionally, AdaSCALE consistently demonstrates superiority in full-spectrum OOD (FSOOD) detection (Yang et al., 2023). Our key contributions are:
@@
We reveal that OOD inputs exhibit more pronounced shifts in top-k activations under minor perturbations compared to ID inputs. Leveraging this, we propose a novel post-hoc OOD detection method using adaptive scaling that attains state-of-the-art OOD detection.
@@
We demonstrate state-of-the-art generalization of AdaSCALE via extensive evaluations across 10 architectures and 3 datasets by tuning mere one hyperparameter for a given setup.
@@ 2 Related Works
Post-hoc methods. Early research on OOD detection primarily focused on designing scoring functions based on logit information (Hendrycks & Gimpel, 2017; Liang et al., 2018; Liu et al., 2020; Hendrycks et al., 2022a; Liu et al., 2023). While these methods leveraged logit-based scores, alternative approaches have explored gradient-based information, such as GradNorm (Huang et al., 2021), GradOrth (Behpour et al., 2023), GAIA (Chen et al., 2023), and Greg-OOD (Sharifi et al., 2025). Given the limited dimensionality of the logit space, which may not encapsulate sufficient information for OOD detection, subsequent studies have investigated activation-space-based methods. These approaches exploit the high-dimensional activations, leading to both parametric techniques such as MDS (Lee et al., 2018), MDS Ensemble (Lee et al., 2018), and RMDS (Ren et al., 2021), as well as non-parametric methods such as KNN-based OOD detection (Sun et al., 2022; Park et al., 2023). Recent advancements have proposed hybrid methodologies that integrate parametric and non-parametric techniques to improve robustness. For instance, ComboOOD (Rajasekaran et al., 2024) combines these paradigms to enhance near-OOD detection performance. Similarly, VIM (Wang et al., 2022) employs a combination of logit-based and distance-based metrics. However, reliance of such approaches on ID statistics (Sun & Li, 2022; Olber et al., 2023; Zhang et al., 2023c) can become a constraint, hindering scalability and practical deployment in real-world applications. To mitigate computational challenges for real-world deployment, recent methods, such as FDBD (Liu & Qin, 2024) and NCI (Liu & Qin, 2023), have focused on enhancing efficiency. Recent advances, such as NECO (Ammar et al., 2024) examines connections to neural collapse phenomena, while WeiPer (Granz et al., 2024), explore class-direction perturbations. Unlike WeiPer, our work deals with perturbation in the input image similar to ODIN (Liang et al., 2018).
@@
Activation-shaping post-hoc methods. A seminal work in OOD detection, ReAct (Sun et al., 2021), identified abnormally high activation patterns in OOD samples and proposed clipping extreme activations. This approach has been further generalized by BFAct (Kong & Li, 2022) and VRA (Xu et al., 2023), which extend activation clipping for enhanced effectiveness. Additionally, BATS (Zhu et al., 2022) refines activation distributions by aligning them with their respective typical sets, while LAPS (He et al., 2024) enhances this strategy by incorporating channel-aware typical sets. Inspired by activation clipping, another line of research explores activation “scaling" as a means to improve OOD detection. ASH (Djurisic et al., 2023) introduces a method to compute a scaling factor as a function of the activation itself, pruning and rescaling activations to enhance the separation of energy scores between ID and OOD samples. However, this approach results in a slight degradation in ID classification accuracy. In response, SCALE (Xu et al., 2024) observes that pruning adversely affects performance and thus eliminates it, leading to improved OOD detection while preserving ID accuracy. SCALE currently represents the state-of-the-art method for ResNet-50-based OOD detection. Despite their efficacy, these activation-based methods exhibit limited generalization across diverse architectures. To address this issue, LTS (Djurisic et al., 2024) extends SCALE by computing scaling factors using post-ReLU activations and applying them directly to logits rather than activations. Our work builds on this line of work, introducing the adaptive scaling mechanism. ATS (Krumpl et al., 2024) argues that relying solely on final-layer activations may result in the loss of critical information beneficial for OOD detection and proposes to leverage intermediate-layer activations too. However, its efficacy is contingent upon the availability of a large number of training samples, whereas our approach attains state-of-the-art performance while utilizing a minimal number of ID samples. A newly proposed method OptFS (Zhao et al., 2024) introduces a piecewise constant shaping function with the goal of generalization across diverse architectures in large-scale settings, while our work exhibits superior generalization extending to small-scale settings too.
@@
Training methods. The training methods incorporate adjustments during training to enhance the ID-OOD differentiating characteristics. They either make architectural adjustments (DeVries & Taylor, 2018; Hendrycks et al., 2019b; Hsu et al., 2020), apply enhanced data augmentations (Xiong et al., 2024; Hendrycks* et al., 2020; Hendrycks et al., 2022b), or make simple training modifications Wei et al. (2022); Regmi et al. (2024a); Zhang et al. (2024). More recent methods have adopted contrastive learning in the context of OOD detection (Ming et al., 2023; Regmi et al., 2024b; Lu et al., 2024; Zou et al., 2025). Moreover, some approaches also either utilize external real outliers (Hendrycks et al., 2019a; Zhang et al., 2023a; Du et al., 2024; Zhu et al., 2023) or synthesize virtual outliers either in image space (Du et al., 2023; Regmi, 2024; Wang et al., 2023; Gao et al., 2023; Zhang et al., 2025; Li et al., 2024; Bai et al., 2024; Nie et al., 2024) or in feature space (Du et al., 2022; Tao et al., 2023; Gao et al., 2024; Li & Zhang, 2025). However, training methods can be costlier and less effective than post-hoc approaches in some large-scale setups (Yang et al., 2022).
@@ 3 Preliminaries
Let denote the input space and denote the label space, where is the number of classes. We consider a multi-class classification setting where a classifier is trained on ID data drawn from an underlying joint distribution , where and . The ID training dataset is denoted as , where is the number of training samples and . The classifier is composed of a feature extractor , and a classifier . The feature extractor maps an input to a feature vector , where and the classifier then maps this feature vector to a logit vector . We refer to individual dimensions of the feature vector as activations, denoted by for the -th dimension. The classifier is trained on to minimize the empirical risk: where is a loss function, such as cross-entropy loss. During inference, the model may encounter data points drawn from a different distribution, denoted as , which is referred to as OOD data. The OOD detection problem aims to identify whether a given input is drawn from marginal distribution or from . Hence, the goal is to design a scoring function that assigns a scalar score to each input , reflecting its likelihood of being an OOD sample. A higher score typically indicates a higher probability of the input being OOD. A threshold is used to classify an input as either ID or OOD: .
@@ 4 Method
In this section, we introduce AdaSCALE, a novel post-processing approach that dynamically adapts the scaling mechanism based on each sample’s estimated OODness. We first revisit and analyze the core principle underlying recent scaling-based static state-of-the-art approaches. Then, we present our key empirical observations regarding activation behavior under minor perturbations, building upon insights from ReAct (Sun et al., 2021). Finally, we detail our proposed adaptive scaling mechanism that leverages these observations to achieve superior OOD detection performance.
4.1 Revisiting Static Scaling Mechanism
Scaling baselines (Djurisic et al., 2023; Xu et al., 2024; Djurisic et al., 2024) use energy score on (directly or indirectly) scaled logits, with higher (magnitude) values indicating higher IDness. They operate by scaling activations / logits with scaling factor computed as:
where denotes the percentile of all elements in activation . While this approach yields sample-specific scaling factors, it imposes a critical constraint: the percentile threshold is static and identical across all test samples, regardless of the nature of samples. We argue that this static nature limits the effectiveness of the scaling procedure and prevents optimal ID-OOD separability. We hypothesize that a proxy for estimating the OODness can be useful for computing dynamic percentiles giving rise to dynamic scaling factors.
@@ 4.2 Observations in Activation Space
A seminal work ReAct (Sun et al., 2021) demonstrated that OOD samples often induce abnormally high activations within neural networks. We extend this finding with an important observation: the positions of such high activations in OOD samples are relatively unstable under minor perturbations compared to ID samples. This instability provides a valuable signal for distinguishing OOD samples from ID samples. Below, we formalize this observation.
@@ 4.2.1 Perturbation Mechanism
Let be an input image with input channels, height, and width. We denote channel value at position as . To identify channel values for perturbation, we employ pixel attribution that quantifies each input element’s influence on the model’s prediction. An attribution function, , assigns a score to each channel value, with lower absolute scores indicating less influence. We select of channel value indices with lowest absolute attribution scores, forming the set . We use a gradient-based attribution , where is predicted class index. To create a perturbed input, we select a subset containing of channel values to perturb. If is perturbation magnitude, the perturbed image is obtained as:
@@ 4.2.2 Activation Shift as OOD Indicator
After obtaining perturbed input , we compute its activation . We define activation shift as the absolute element-wise difference between original activation and the perturbed one:
@@
Figure 2 illustrates the key insight of our approach: activation shift at extreme (high-magnitude) activations is consistently more pronounced in OOD samples compared to ID samples. This behavior can be understood intuitively: ID samples activate network features in a stable, predictable manner reflecting learned patterns, while OOD samples trigger less stable, more arbitrary high activations that shift significantly under perturbation. Based on this observation, we propose using activation shift at the top- highest activations as a metric to estimate OODness of a sample:
@@
where returns the indices of the highest values in . As evidenced by ratio () shown in Figure 3, the statistic generally assigns higher values to OOD samples than ID ones. However, the high variance of metric (Figure˜2) suggests the possibility of overoptimistic estimations. To address this issue, we introduce a correction term that exhibits an opposing behavior: it tends to be higher for ID samples than for OOD samples. Figure 5 in Appendix˜C shows that the perturbed activations of ID samples tend to be higher than those of OOD ones, especially in high-activation regions. We leverage this complementary signal by defining , where is a hyperparameter denoting the number of considered activations. We refine our OOD quantification by combining both metrics, weighted by a hyperparameter :
@@
Indeed, Figure˜3 illustrates that , suggesting that the correction term helps mitigate overconfident estimations. If be the set of values on ID validation samples, we could transform any into a normalized probability scale by constructing empirical cumulative distribution function (eCDF) derived from . The eCDF, denoted as , can be defined as: , where is the indicator function. A higher value of indicates a higher likelihood of the sample being OOD. Importantly, our experiments suggest that as few as 10 ID validation samples are sufficient to construct an effective eCDF for this purpose (See Table˜9).
4.3 Proposed Approach: Adaptive Scaling
Building on our observations, we propose AdaSCALE (Adaptive SCALE), a novel approach that introduces dynamic, sample-specific adjustments to the scaling procedure. The key insight is that percentile threshold should be a function of each test sample’s estimated OODness rather than a fixed value. The scaling factor increases as the percentile threshold rises (i.e., when more activations are excluded from the denominator in Equation 1). For optimal ID-OOD separation, we must scale ID samples more strongly than OOD samples, requiring a higher percentile for ID samples. We define an adaptive percentile threshold as: , where and are hyperparameters that define the minimum and maximum limits of percentile threshold. It ensures samples with lower estimated OODness receive higher percentile thresholds, resulting in stronger scaling. (See Algorithm˜1). We implement two variants: AdaSCALE-A scales activations as (Djurisic et al., 2023; Xu et al., 2024). AdaSCALE-L scales logits as (Djurisic et al., 2024). This approach is also outlined in Figure˜4.
5 Experiments
We use pre-trained models provided by PyTorch for ImageNet-1k experiments. For CIFAR experiments, we train three models per network using the standard cross-entropy loss and report the mean results across these three independent trials. The evaluation setup is provided in Table˜1. Although we tune for optimal results in each case (see Appendix˜G), near-optimal performance can be achieved by tuning only while fixing to across all 22 cases (See Table˜5). These values were determined via automatic parameter search (Yang et al., 2022; Zhang et al., 2023b) with respect to the AUROC metric using ResNet-50 model on ImageNet-1k benchmark. Best results are bold, and second-best results are underlined.
Conventional OOD detection ID datasets Near-OOD Far-OOD Network CIFAR-10/100 CIFAR-100/10 MNIST, SVHN, WRN-28-10, DenseNet-101 TIN Textures, Places365 ImageNet-1k SSB-Hard iNaturalist, EfficientNetV2-L, ResNet-101 NINCO OpenImage-O DenseNet-201, ViT-B-16 ImageNet-O Textures ResNet-50, ResNeXt-50 Places RegNet-Y-16, Swin-B Covariate shifted datasets for full spectrum OOD detection ImageNet-1k ImageNet-C, ImageNet-R, ImageNet-V2, ImageNet-ES
Baselines. MSP (Hendrycks & Gimpel, 2017), EBO (Liu et al., 2020), ReAct (Sun et al., 2021), MLS (Hendrycks et al., 2022a), ASH (Djurisic et al., 2023), SCALE (Xu et al., 2024), BFAct (Kong & Li, 2022), LTS (Djurisic et al., 2024), OptFS (Zhao et al., 2024). Currently, SCALE is the best-performing method (with ResNet-50), while OptFS is the best-generalizing method.
5.1 Empirical Results